OpenStudy (anonymous):

Find the coefficient of the squared term in the simplified form for the second derivative, f "(x) for f(x) = (x3 + 3x2 + 3)(3x3 - 6x2 - 8x + 1) . Use the hyphen symbol, -, for negative values

OpenStudy (anonymous):

We have to find the coefficient of squared term of the second derivative of the function. When f(x)=x^4, f''(x)= 12x^2 =>12 is the coefficient. Now we need only the term in degree 4 in the given function. x^3*-8x +3x^2*-6x^2 = -42x^4 => the required coefficient is 12*-42 = -504